robinson crusoe produces upper-loop product of $1000. he pays $750 in wages, $125 in interest, and $75 in rent. what must his profit be? if three-fourths of crusoe’s output is consumed and the rest invested, calculate crusoeland’s gdp with both the product and the income approaches and show that they must agree exactly.
1
Expert's answer
2021-08-24T10:17:54-0400

Profit is given by total revenue less all expenses(total cost) incurred

robinson crusoe revenus is $ 1000

Expenses of robinson crusoe

Rent $ 75

Interest $ 125

Wages $ 750

Total expenses = $ 950

Profit = 1000 - 950

Profit = $ 50


GDP

Product approach

GDP = C + I

Consumption is THREE QUARTERS OF 1000 = 750

Investment = 1000- consumption

Investment= 1000- 750

Investment = 250

Therefore GDP = 750 + 250

GDP= $1000


INCOME APPROACH

GDP = Wages + Rent + Interest + Profit

GDP = 750 + 75 + 125 + 50

GDP = $1000


Conclusion, both product approach and income gives the same GDP of $1000.